# A Comprehensive Survey on Deep Multimodal Learning with Missing Modality

### v2

@@ 3. Applications and Datasets
Over the past decade, multimodal learning based on deep learning has experienced explosive growth, leading to extensive applications in both academia and industry. Accompanying these applications is the emergence of various forms of multimodal datasets. However, the collection of such datasets is often labor-intensive and costly. In certain specific application directions, issues such as user privacy concerns, sensor malfunctions on data collection devices, and other factors can result in datasets with missing modalities. In severe cases, up to 90% of the samples may have missing modalities, making it challenging for conventional MLFM to achieve good performance during model training. This has given rise to the task of MLMM. Since the factors causing incomplete datasets usually stem from different application directions, we below introduce the corresponding datasets based on these applications: Sentiment Analysis, Medical Diagnosis, Retrieval/Captioning, Remote Sensing, Robotic Vision, and others. We also categorize those datasets according to the applications and data type in Table 3.
Table 3. Common and typical datasets used by deep MLMM methods divided by main applications and modality type. “Vision” includes data from visual sensors such as RGB, Depth, Infrared, LiDAR, Radar, Event, and Optical Flow. “Bio-Sensors” includes data from sensors like: electrophysiological and respiratory sensors.
Applications Modality Type Common and Typical Datasets
@@
Medical Diagnosis MRI, CT Scans BRATS (Menze et al., 2014; Myronenko, 2019), ADNI (Jack Jr et al., 2008), IXI [^1] Source: https://brain-development.org/ixi-dataset/
@@
Remote Sensing Vision WHU-OPT-SAR (Li et al., 2022b), DFC2020 (Yokoya et al., 2020), MSAW (Shermeyer et al., 2020), Trento (Rasti et al., 2017), Houston (Piacentini, 2023), Berlin (Hong et al., 2021), GRSS (Debes et al., 2014)
@@ 3.1. Sentiment Analysis
Typically, the goal of sentiment analysis is to classify the current emotional state by combining information from multiple modalities such as textual, auditory, and RGB information. This field has garnered significant attention from researchers due to its promising applications in various specific scenarios, including customer service, market research, health monitoring, entertainment, and advertising. In the early stages of audio-visual sentiment analysis research, researchers discovered that facial detection algorithm in data collection devices sometimes failed to capture faces (e.g., due to occlusion) or the recorded audio is too noisy to use (Cohen et al., 2004; Sebe et al., 2005), resulting in some data samples that only contained audio or image, leading to incomplete audio-visual datasets. Consequently, researchers began exploring methods to enable sentiment analysis models to continue functioning effectively when inputting data samples with missing modalities, and additionally incorporating textual data as an extra modality. Currently, the datasets for applying sentiment analysis can be roughly divided into two categories. One type of datasets  (Martin et al., 2006; Livingstone and Russo, 2018; Cao et al., 2014; Stappen et al., 2021a, b; Ringeval et al., 2013; Zheng and Lu, 2015; Zadeh et al., 2018; Busso et al., 2008, 2016; Poria et al., 2018; Yu et al., 2020; Wöllmer et al., 2013; Morency et al., 2011) involves using video, audio, text, and biosensors to determine human emotional states in real life or in movies. The other type (Guérin et al., 2013; Nguyen et al., 2018; Matsui et al., 2017) involves using text and images to assess emotional states in contexts such as comic books. Please refer to Table 3 for the classification based on modality type.
@@ 3.2. Medical Diagnosis
Medical diagnosis has always been the focus of attention of people all over the world, but the uneven distribution of medical resources, how to promote personalized medicine, improve medical efficiency and how to reduce diagnostic errors are all current pain points. The emergence of artificial intelligence is expected to improve these problems. Because medical diagnosis requires comprehensive judgment of information from various modalities such as medical history, physical examination and imaging data, which is exactly what multimodal learning is good at, most researchers are currently developing intelligent medical diagnosis systems based on multimodal learning. The datasets that current attempts used to solve the missing modality problem in multimodal medical diagnosis mainly focus on eight areas, such as Neuroimaging and Brain Disorders (Menze et al., 2014; Myronenko, 2019; Young et al., 2017; Jack Jr et al., 2008; Lesjak et al., 2018), Cardiovascular (Liu et al., 2016; Moody and Mark, 2001), Cancer Detection (Arya and Saha, 2020; Tomczak et al., 2015), Women Health Analysis (Zhang et al., 2023c; Xu et al., 2021), Ophthalmology (Li et al., 2021; Zhang et al., 2022), Sleep Disorders (Lee et al., 2022; Marcus et al., 2013), Clinical Predictions (Johnson et al., 2016, 2023, 2019), and Biomedical Analysis (Chaptoukaev et al., 2024; Schmidt et al., 2018; Kemp et al., 2000). We have categorized some commonly used and representative datasets from above by modality type, as shown in Table 3.
@@ 3.3. Information Retrieval
Information Retrieval is a technology that automatically retrieves relevant content or data based on queries, historical behavior, preferences, and attribute data. It has received widespread attention due to its success on various platforms such as search engines and social media. It improves user satisfaction and information access experience through algorithmic analysis and prediction of content that users may find relevant. Multimodal learning makes modern retrieval systems possible because it can effectively process and analyze multiple types of information including text, images, audio, etc. Due to user privacy concerns and data sparsity issues, some researchers have begun exploring MLMM approaches in retrieval systems. Many of these studies leverage datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017; Duin, [n. d.]; Chua et al., 2009; Xu et al., 2016; Lin et al., 2023) from websites such as Amazon and TikTok for their research. See categories of those datasets based on modality type in Table 3.
@@ 3.4. Remote Sensing
By leveraging multimodal learning, we can integrate various types of visual data, such as Synthetic Aperture Radar data and multi-/hyper-spectral data. As a result, corresponding various datasets have also emerged (Yokoya et al., 2020; Shermeyer et al., 2020; Rasti et al., 2017; Piacentini, 2023; Hong et al., 2021; Debes et al., 2014). Those dataset enables the assessment and analysis of different environmental conditions, resources, and disasters on earth from satellites or aircraft. Such capabilities can significantly aid in environmental protection, resource management, and disaster response. In practical, in remote sensing tasks such as multimodal land cover classification tasks, optical modalities may be unavailable due to cloud cover or sensor damage. Therefore, it is also very necessary to solve the missing modality problem for remote sensing researchers.
@@ 3.5. Robotic Vision
Robotic vision is the technology that enables robots to perceive and understand their environment through visual sensors by acquiring and processing image data. This includes tasks such as object and facial recognition, environment modeling, navigation, and behavior understanding, allowing robots to autonomously perform complex operations and interact with humans. Typically, the modality combination formula in robotic vision tasks can be expressed as RGB+X, where X can include, but are not limited to, LiDAR, radar, infrared sensors, depth sensors and auditory sensors. Both RGB and other sensor data can be missing for various reasons. Below, we introduce five common tasks in robotic vision concerning the missing modality problems. (1) Multimodal Segmentation: This task aims to input multimodal data and use segmentation masks to locate objects of interest. In the context of missing modalities, this task is common in datasets for autonomous vehicles, where sensors might fail due to damage or adverse weather conditions. Typical data combinations include RGB+(Depth/Optical Flow/LiDAR/Radar/Infrared/Event data) (Takumi et al., 2017; Liu et al., 2023b; Caesar et al., 2020; Zhang et al., 2023a). Other datasets focus on indoor scene segmentation (Silberman et al., 2012; Ha et al., 2017) (RGB, Depth, Thermal) and material segmentation (Liang et al., 2022) (RGB, Angle of Linear Polarization, Degree of Linear Polarization, Near-Infrared images). (2) Multimodal Detection: Similar to multimodal segmentation, multimodal detection aims to locate objects of interest using bounding boxes. Commonly used multimodal detection combinations in the context of missing modalities include: RGB with Depth (Silberman et al., 2012; Lai et al., 2011) and RGB with Thermal (Hwang et al., 2015; kaggle, 2020). (3) Multimodal Activity Recognition: This task can input data from the aforementioned visual sensors and also use audio cues to recognize human activities. Common datasets (Grauman et al., 2022; Huh et al., 2023; Damen et al., 2022; Gong et al., 2023; Wang et al., 2014; Shahroudy et al., 2016) for this task are combinations of RGB with Audio, Depth, Thermal, Inertial and skeleton data. (4) Multimodal Person Re-Identification: This task aims to use depth and thermal (Ye et al., 2021; Wu et al., 2017) information to robustly identify individuals under varying lighting conditions, such as at night. (5) Multimodal Face Anti-Spoofing: This task utilizes multiple sensors (such as RGB, infrared, and depth cameras) and multiple types of information to detect and prevent spoofing in facial recognition systems (John and Kawanishi, 2022; Zhang et al., 2020; Liu et al., 2021b). Its purpose is to enhance the system’s ability to identify fake faces (such as photos, videos, masks, etc.) by comprehensively analyzing data from different modalities, thereby increasing the security and reliability of facial recognition systems.
3.6. Other Applications
@@
Except above tasks, there are many other areas where MLMM methods are being explored. For example, conventional audio-visual classification (Vielzeuf et al., 2018); audio-visual question answering (Park et al., 2024); multimodal large language models in visual-dialogue (Bubeck et al., 2023), captioning (Liu et al., 2023a) and so on; hand pose estimation using depth images and heat distributions (Tompson et al., 2014; Choi et al., 2017); knowledge graph completion utilizing multimodal data with missing modalities (Toutanova et al., 2015; Liu et al., 2019); multimodal time series prediction, such as stock prediction (Luo et al., 2024) and air quality forecasting (Zheng et al., 2015); multimodal gesture generation on how to create natural animations (Liu et al., 2022); analyzing single-cell multimodal data (Mimitou et al., 2021; Swanson et al., 2021) in biology.
3.7. Discussion
@@
Current MLMM research primarily focuses on the areas of sentiment analysis, medical diagnosis, remote sensing analysis, retrieval/captioning, and robotic vision. Among these, a significant portion of research is dedicated to addressing the missing modality problem in video, text, and audio-based sentiment analysis, MRI segmentation and clinical prediction in medical diagnosis, and multi-sensor semantic segmentation for autonomous vehicles. In contrast, research efforts in MLMM for streaming data and scientific fields are relatively limited. In addition, the majority of the publicly available datasets mentioned above are complete in terms of modalities, and naturally occurring datasets with missing modalities are rare. As a result, these tasks are often evaluated by considering all possible combinations of missing modalities based on the existing types of modalities in the dataset, and then performing training and testing on the performance of the missing modalities, supplemented by different missing modality rates. In the works we reviewed, 38% controlled for varying degrees of missing modality rates and validated accordingly, while the remaining 62% used random modality missing methods. When categorizing missing modality rates into mild (¡30%), moderate (30%-70%), and severe (¿70%) levels, the proportions are 72.6%, 82.2%, and 69.9%, respectively. It is worth noting that a single study may perform validation across multiple missing modality rate levels. Additionally, we found that 36.4% of the works trained on incomplete training datasets (cannot access missing modality data during training), while 63.6% actually focused on testing with missing modalities. We can see that there is still small number of research on incomplete training datasets and call on more researchers to pay attention to this field.
@@ 4. Open Issues and Future Research Directions
4.1. Debate Between Recovery and Non-Recovery Methods
@@
In MLMM, “to recover or not to recover the missing modality”, that is also the question. According to our observations, most current approaches to handling MLMM can be broadly categorized into two types: one focuses on recovering the missing modal information so that the multimodal model can continue functioning (Wang et al., 2023a; Ma et al., 2021a), and the other attempts to make predictions using only the currently available modal information (Hu et al., 2020; Zhang et al., 2024). However, some studies (Yao et al., 2024) have shown that the recovered modal information might be ignored or dominated by the existing modal information due to the imbalance missing rates of different modalities in the training dataset, leading the model to still rely on the available modalities when making predictions. Moreover, other research (Lin et al., 2024) has found that in situations where crucial modal information is missing, the model might be unduly influenced by less important existing modal information, resulting in incorrect outcomes. In such cases, attempting to recover missing modal information can potentially alleviate this problem. Consequently, determining under what circumstances recovery or non-recovery methods are more beneficial for model predictions, as well as how to measure whether the recovered modal information is not dominated by other modal information or is unbiased, remains significant challenges. Addressing these issues is essential for advancing the effectiveness and reliability of deep multimodal learning with missing modality methods.
4.2. Efficiency and Lightweight
Current solutions for improving the performance of MLMM often overlook the need for more lightweight and efficient methods. For instance, some model selection methods (Hu et al., 2020; Xue and Marculescu, 2023) require training an independent model for each potential cases of missing modalities or a combination of many uni- and multi-modal models. Similarly, methods (Chen et al., 2024b; Zhang et al., 2023b) that aim to recover missing modal information typically involve using a distinct model for each kind of missing information or a large unified model to generate all modalities. Although these approaches generally perform well, they are often too heavy. In the real world, many multimodal models need to be deployed on resource-constrained devices, such as space or disaster-response robots, and mobile devices. These devices cannot accommodate high-performance GPUs and are prone to sensor damage, which can be difficult or costly to repair. Therefore, there is an urgent need to explore efficient and lightweight solutions for MLMM that can operate effectively under these constraints. Developing such solutions is critical for the practical deployment of MLMM models in real-world where resources are limited.
4.3. Multimodal Streaming Data
Currently, only part of research focuses on handling missing modalities in multimodal streaming data, mainly in sentiment analysis (Lin and Hu, 2023). However, long sequences of multimodal data, such as RGB+X video and multimodal time series data, are common in the real world and are essential for tasks like anomaly detection and video tracking on robots. Therefore, MLMM needs to address missing modality problem in streaming data to advance this field further.
4.4. Multimodal Reinforcement Learning
Multimodal reinforcement learning leverages information from different sensory modalities, enabling agents to learn effective strategies in environments. This approach has wide applications in areas such as robotic grasping, drone control, and driving decisions. However, these agent-based tasks often face practical challenges, as real-world scenarios frequently encounter sensor failures or restricted access to certain modal data, compromising the robustness of the algorithms. Currently, only a limited amount of work (Vasco et al., 2021; Mei et al., 2023; Lee et al., 2021; Chen et al., 2021; Awan et al., 2022) try to address the problem of missing data or modalities in reinforcement learning. Therefore, we hope the community can focus more on this practically significant missing modality problem in the reinforcement learning.
4.5. Multimodal AI for Natural Science
In scientific domains such as drug prediction (Deng et al., 2020), materials science (Muroga et al., 2023), and astronomy (Hong et al., 2023), the application of multimodal learning is of paramount importance. These fields inherently involve diverse types of data, from molecular structures and genomic sequences to spectral images and material properties. Multimodal learning holds the promise of integrating these varied data sources to enhance predictive accuracy and uncover novel insights. However, the practical implementation of multimodal learning in scientific research faces significant challenges. Data access restrictions, high acquisition costs, and the incompatibility of heterogeneous data sources often result in incomplete datasets or entirely missing modalities. These issues hinder the seamless integration of multimodal data, making it difficult to fully leverage the potential of multimodal learning in scientific experiments. Despite the critical need, there has been limited research (He et al., 2024) focusing on MLMM within the scientific field. Addressing this gap is crucial for advancing AI-driven scientific discoveries. Innovative approaches are required to handle incomplete multimodal dataset and develop robust models capable of performing well even when some modalities are absent. This will not only improve the applicability of AI in scientific research but also pave the way for breakthroughs in understanding and innovation.
4.6. Lack of Multimodal Large Language Model Benchmarks

### v3

@@ 6. Applications and Datasets
The collection of multimodal datasets is often labor-intensive and costly. In certain specific application directions, issues such as user privacy concerns, sensor malfunctions on data collection devices, and other factors can result in datasets with missing modalities. In severe cases, up to 90% of the samples may have missing modalities, making it challenging for conventional MLFM to achieve good performance. This has given rise to the task of MLMM. Since the factors causing incomplete datasets usually stem from different application directions, we introduce the following datasets based on the applications which is common in MLMM tasks: Sentiment Analysis, Medical Diagnosis, Retrieval/Captioning, Remote Sensing, Robotic Vision, and others. We categorize those datasets according to the applications and data type in Table 3.
Table 3. Common datasets used by deep MLMM methods divided by main applications and modality types. “Vision” includes data from visual sensors such as RGB images, Depth, Infrared, LiDAR, Radar, Event, and Optical Flow. “Bio-Sensors” includes data from sensors like: electrophysiological and respiratory sensors. Other modalities include audio, text, CT scans, MRI, and skeleton.
Applications Modality Types Common and Typical Datasets
@@
Medical Diagnosis MRI, CT Scans BRATS-series (Menze et al., 2014; Myronenko, 2019), ADNI (Jack Jr et al., 2008), IXI [https://brain-development.org/ixi-dataset]
@@
Remote Sensing Vision WHU-OPT-SAR (Li et al., 2022c), DFC2020 (Yokoya et al., 2020), MSAW (Shermeyer et al., 2020), Trento (Rasti et al., 2017), Houston (Piacentini, 2023), Berlin (Hong et al., 2021), GRSS (Debes et al., 2014)
@@ 6.1. Sentiment Analysis
Typically, the goal of sentiment analysis is to classify the current emotional state by combining information from multiple modalities such as textual, auditory, and RGB information. This field has garnered significant attention due to its promising applications in various specific scenarios, including market research, health monitoring, and advertising. In the early stages of audio-visual sentiment analysis research, researchers discovered that facial detection algorithm in data collection devices sometimes failed to capture faces (e.g., due to occlusion) or the recorded audio is too noisy to use (Cohen et al., 2004; Sebe et al., 2005), resulting in some samples that only contained audio or image. Consequently, researchers began exploring methods to enable sentiment analysis models to continue functioning effectively when facing missing-modality samples. Currently, the datasets for applying sentiment analysis can be roughly divided into two categories. One type of dataset  (Martin et al., 2006; Livingstone and Russo, 2018; Cao et al., 2014; Stappen et al., 2021a, b; Ringeval et al., 2013; Zheng and Lu, 2015; Zadeh et al., 2018; Busso et al., 2008, 2016; Poria et al., 2018; Yu et al., 2020; Wöllmer et al., 2013; Morency et al., 2011) involves using video, audio, text, and biosensors to determine human emotional states in real life or in movies. The other type (Guérin et al., 2013; Nguyen et al., 2018; Matsui et al., 2017) involves using text and images to assess emotional states in contexts such as comic books. Please refer to Table 3 for details.
@@ 6.2. Medical Diagnosis
Medical diagnosis requires comprehensive judgments from various modalities such as medical history, physical examination and imaging data, which is exactly what multimodal learning is good at, many researchers developed multimodal intelligent medical diagnosis systems. The current datasets for multimodal medical diagnosis that contain data with missing modality problems are focused on on eight areas: Neuroimaging and Brain Disorders (Menze et al., 2014; Myronenko, 2019; Young et al., 2017; Jack Jr et al., 2008; Lesjak et al., 2018), Cardiovascular (Liu et al., 2016; Moody and Mark, 2001), Cancer Detection (Arya and Saha, 2020; Tomczak et al., 2015), Women Health Analysis (Zhang et al., 2023c; Xu et al., 2021), Ophthalmology (Li et al., 2021; Zhang et al., 2022), Sleep Disorders (Lee et al., 2022; Marcus et al., 2013), Clinical Predictions (Johnson et al., 2016, 2023, 2019), and Biomedical Analysis (Chaptoukaev et al., 2024; Schmidt et al., 2018; Kemp et al., 2000). We have categorized some commonly used and representative datasets from list above by modality type, as shown in Table 3.
@@ 6.3. Information Retrieval
Information retrieval is a technology that automatically retrieves relevant content or data based on queries, historical behavior, preferences, and attribute data. It has received widespread attention due to its success on various platforms such as search engines and social media. It improves user satisfaction and information access experience through algorithmic analysis and prediction of content that users may find relevant. Multimodal learning makes modern retrieval systems possible because it can effectively process and analyze multiple types of information, including text, images, audio, etc. Due to user privacy concerns and data sparsity issues, some researchers have begun exploring MLMM approaches in retrieval systems. Many of these studies leverage datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017; Duin, [n. d.]; Chua et al., 2009; Xu et al., 2016; Lin et al., 2023) from websites such as Amazon and TikTok for their research. We list those datasets based on modality type in Table 3.
@@ 6.4. Remote Sensing
By leveraging multimodal learning, we can integrate various types of visual data, such as Synthetic Aperture Radar data and multi-/hyper-spectral data. As a result, various datasets have emerged (Yokoya et al., 2020; Shermeyer et al., 2020; Rasti et al., 2017; Piacentini, 2023; Hong et al., 2021; Debes et al., 2014). Those datasets enable the assessment and analysis of different environmental conditions, resources, and disasters on earth from satellites or aircraft. Such capabilities can significantly aid in environmental protection, resource management, and disaster response. In practice, remote sensing tasks, such as multimodal land cover classification tasks, optical modalities may be unavailable due to cloud cover or sensor damage. Therefore, it is also necessary to address the missing modality problem for remote sensing problems.
@@ 6.5. Robotic Vision
Robotic vision is the field that enables robots to perceive and understand their environment through visual sensors by acquiring and processing image data. This includes tasks such as object and facial recognition, environment modeling, navigation, and behavior understanding, allowing robots to autonomously perform complex operations and interact with humans. Typically, the modality combination formula in robotic vision tasks can be expressed as RGB+X, where X can include, but are not limited to, LiDAR, radar, infrared sensors, depth sensors and auditory sensors. Both RGB and other sensor data can be missing for various reasons. Below, we introduce five common tasks in robotic vision concerning the missing modality problems. (1) Multimodal Segmentation: This task aims to input multimodal data and use segmentation masks to locate objects of interest. In the context of missing modalities, this task is common in datasets for autonomous vehicles, where sensors might fail due to damage or adverse weather conditions. Typical data combinations include RGB+(Depth/Optical Flow/LiDAR/Radar/Infrared/Event data) (Takumi et al., 2017; Liu et al., 2023b; Caesar et al., 2020; Zhang et al., 2023a). Other datasets focus on indoor scene segmentation (Silberman et al., 2012; Ha et al., 2017) (RGB, Depth, Thermal) and material segmentation (Liang et al., 2022) (RGB, Angle of Linear Polarization, Degree of Linear Polarization, Near-Infrared images). (2) Multimodal Detection: Similar to multimodal segmentation, multimodal detection aims to locate objects of interest using bounding boxes. Commonly used multimodal detection combinations in the context of missing modalities include: RGB with Depth (Silberman et al., 2012; Lai et al., 2011) and RGB with Thermal (Hwang et al., 2015; kaggle, 2020). (3) Multimodal Activity Recognition: This task can input data from the aforementioned visual sensors and also use audio cues to recognize human activities. Common datasets (Grauman et al., 2022; Huh et al., 2023; Damen et al., 2022; Gong et al., 2023; Wang et al., 2014; Shahroudy et al., 2016) for this task are combinations of RGB with Audio, Depth, Thermal, Inertial and skeleton data. (4) Multimodal Person Re-Identification: This task aims to use depth and thermal (Ye et al., 2021; Wu et al., 2017) information to robustly identify individuals under varying lighting conditions, such as at night. (5) Multimodal Face Anti-Spoofing: This task utilizes multiple sensors (such as RGB, infrared, and depth cameras) to detect and prevent spoofing in facial recognition systems (John and Kawanishi, 2022; Zhang et al., 2020; Liu et al., 2021b). Its purpose is to enhance the system’s ability to identify fake faces (such as photos, videos, masks, etc.) by comprehensively analyzing different modalities, thereby increasing the security and reliability of facial recognition systems.
@@
Other Applications: There are many other areas where MLMM methods are being explored. For example, conventional audio-visual classification (Vielzeuf et al., 2018); audio-visual question answering (Park et al., 2024); multimodal large language models in visual-dialogue (Bubeck et al., 2023), captioning (Liu et al., 2023a); hand pose estimation using depth images and heat distributions (Tompson et al., 2014; Choi et al., 2017); knowledge graph completion utilizing multimodal data with missing modalities (Liu et al., 2019); multimodal time series prediction, such as stock prediction (Luo et al., 2024) and air quality forecasting (Zheng et al., 2015); multimodal gesture generation on how to create natural animations (Liu et al., 2022); multimodal analysis of single-cell data (Mimitou et al., 2021) in biology.
@@
Discussion: Current MLMM research primarily focuses on the areas of sentiment analysis, medical diagnosis, remote sensing analysis, retrieval/captioning, and robotic vision. Among these, a significant portion of research is dedicated to addressing the missing modality problem in video, text, and audio-based sentiment analysis, MRI segmentation and clinical prediction in medical diagnosis, and multi-sensor semantic segmentation for autonomous vehicles. In contrast, research efforts in MLMM for streaming data and scientific fields are relatively limited. In addition, the majority of the publicly available datasets mentioned above are complete in terms of modalities, and naturally occurring datasets with missing modalities are rare. As a result, these tasks are often evaluated by considering all possible combinations of missing modalities based on the existing types of modalities in the dataset, and then performing training and testing on the performance of the missing modalities, supplemented by different missing modality rates. In the works we reviewed, 38% controlled for varying degrees of missing modality rates and tested accordingly, while the remaining 62% used random modality missing rates for training and testing. When categorizing missing modality rates into mild (¡30%), moderate (30%-70%), and severe (¿70%) levels, the proportions are 72.6%, 82.2%, and 69.9%, respectively. It is worth noting that a single study may perform validation across multiple missing modality rate levels. Additionally, we found that 36.4% of the works trained on incomplete training datasets (cannot access missing modality data during training), while 63.6% actually focused on testing with missing modalities. We can see that there is still small number of research on incomplete training datasets, so it is important to emphasise the need for more research to this field.
@@ 7. Open Issues and Future Research Directions
7.1. Accurate Missing Modality/Representation Data Generation
Many works (Ma et al., 2021b) have shown that recovering, reconstructing or generating missing modalities or their representations can help improve the performance of the model when encountering missing-modality samples. However, these generated modalities or representations usually have artifacts or hallucinations, which may be caused by the generative model itself or its training datasets. Therefore, exploring how to generate accurate and unbiased missing modalities or their representations will be the important research direction in the future.
7.2. Recovery or Non-Recovery Methods ?
@@
In MLMM, “to recover or not to recover the missing modality”, that is the question. According to our observations, most current approaches to handling MLMM can be broadly categorized into two types: one focusing on recovering the missing modal information so that the multimodal model can continue functioning (Wang et al., 2023a; Ma et al., 2021b), and the other attempting to make predictions using only the data modalities available (Hu et al., 2020; Zhang et al., 2024). However, some studies (Yao et al., 2024) have shown that the recovered modality information might be ignored or dominated by the existing modality information due to the imbalance missing rates of different modalities in the training set, leading the model to still rely on the available modalities when making predictions. Moreover, Lin et al. (Lin et al., 2024) have found that in situations where crucial modality information is missing, the model might be unduly influenced by less important existing modality information, resulting in incorrect outcomes. In such cases, attempting to recover missing modality information can potentially alleviate this problem. Consequently, determining under what circumstances recovery or non-recovery methods are more beneficial for model predictions, as well as how to measure whether the recovered modality information is not dominated by other modality information or is unbiased, remains a significant challenge. Addressing these issues is essential for advancing the effectiveness and reliability of deep multimodal learning with missing modality methods.
7.3. Benchmarking and Evaluations for Missing Modality Problems